Three pipes deliver water to a large exit pipe. The velocity at section 2 is 5 ms-1, and the exit flow rate, Q4 is 120 m3hr-1. Find the velocities at 1, 3 and 4 if it is known that Q3/Q4 = 0.5



1
Expert's answer
2021-06-14T17:04:02-0400

Continuity equation for fixed CV with ID flow and discrete CS's

"0= \sum m_ {out}- \sum m_{in}"

where m "= \rho Q= \rho VA" , thus

"0=(\rho Q_4)-( \rho V_1A_1+\rho V_2A_2+\rho V_3A_3)"

Since "Q_3=0.5Q_4"

"V_3= \frac{Q_4}{0.5A_3}= \frac{120}{\frac{\pi}{2}0.06^2}=5.89 m/hr"

substituting "V_3"

"0=120 -(V_1 \frac{\pi}{4}0.04^2+5 \frac{\pi}4 0.05^2+5.98 \frac{\pi}40.06^2)"

"V_1=5.43 m\/hr"